# Low solution rank of the matrix LASSO under RIP with consequences for rank-constrained algorithms

### v1

@@ Abstract
We show that solutions to the popular convex matrix LASSO problem (nuclear-norm–penalized linear least-squares) have low rank under similar assumptions as required by classical low-rank matrix sensing error bounds. Although the purpose of the nuclear norm penalty is to promote low solution rank, a proof has not yet (to our knowledge) been provided outside very specific circumstances. Furthermore, we show that this result has significant theoretical consequences for nonconvex rank-constrained optimization approaches. Specifically, we show that if (a) the ground truth matrix has low rank, (b) the (linear) measurement operator has the matrix restricted isometry property (RIP), and (c) the measurement error is small enough relative to the nuclear norm penalty, then the (unique) LASSO solution has rank (approximately) bounded by that of the ground truth. From this, we show (a) that a low-rank–projected proximal gradient descent algorithm will converge linearly to the LASSO solution from any initialization, and (b) that the nonconvex landscape of the low-rank Burer-Monteiro–factored problem formulation is benign in the sense that all second-order critical points are globally optimal and yield the LASSO solution.
@@ 1 Introduction and main result
Because we hypothesize that $M_{*}$ is low-rank, we will use the popular nuclear-norm–regularized least-squares estimator:
@@
Our goal in this paper is to show that, under similar conditions to those allowing accurate recovery, we can furthermore guarantee that any solution $\widehat{M}$ to (2) itself has low rank. Although promoting low rank was always the goal of the nuclear norm regularizer, we are not aware of prior theoretical guarantees of this for general measurements $\mathcal{A}$. Low rank of $\widehat{M}$ is useful for multiple reasons:
  • Storage and computation: If $\widehat{M}$ has rank $\hat{r}\ll\min\{d_{1},d_{2}\}$, it can be stored in $O(\hat{r}(d_{1}+d_{2}))$ space compared to $O(d_{1}d_{2})$ space for a general dense matrix. Similarly, the cost of many computations (e.g., matrix multiplication or matrix-vector products) can also be reduced by a factor of approximately $\frac{\hat{r}}{\min\{d_{1},d_{2}\}}$.
  • Estimation algorithm efficiency: The convex program (2) is a semidefinite program with $O(d_{1}d_{2})$ variables. Although this problem can be solved in polynomial time, the computation may still be impractically expensive when $d_{1}$ and $d_{2}$ are very large. Instead, when we know that the solution has low rank (say, bounded by $r$), we can optimize directly over rank-$r$ representations with $O(r(d_{1}+d_{2}))$ variables. This is much cheaper computationally and is very popular in practice. We explore this in much greater detail in Section 2. We consider both a rank-restricted iterative spectral algorithm and a Burer-Monteiro factored representation, and we prove that, under similar conditions as required by our rank bound Theorem 1, these algorithmic approaches are guaranteed to find the global optimum of (2).
Our result will depend on three properties of our measurements in (1):
@@
With these concepts in place, we can state our main result (proved in Section 4):
@@
[…]
Note that there is some flexibility in how we choose $M_{*}$ and $\xi$. For example, if the ground truth is only approximately low-rank, we can decompose it into an exactly low-rank matrix plus a residual and then set $M_{*}$ to be the low-rank matrix and subsume the residual (through $\mathcal{A}$) into $\xi$. Again, see [1, 2] for examples of this in error analysis.
@@
The exact solution can be obtained by singular value soft-thresholding $Y$: if $Y=U_{Y}\Sigma_{Y}V_{Y}^{T}$ is the singular value decomposition (SVD) of $Y$, then the optimal solution is $\widehat{M}=U_{Y}(\Sigma_{Y}-\lambda I)_{+}V_{Y}^{T}$, where $(\Sigma_{Y}-\lambda I)_{+}$ replaces any negative diagonal elements of $\Sigma_{Y}-\lambda I$ by zero. If $\lVert\Xi\rVert_{\ell_{2}}\leq\lambda$, simple singular value perturbation arguments yield $\operatorname{rank}(\widehat{M})\leq\operatorname{rank}(M_{*})$. Our result Theorem 1 can be interpreted as extending this (modulo constants and the RIP constant $\delta^{*}$) to general RIP measurement operators $\mathcal{A}$.
@@ 2 Guarantees for nonconvex rank-constrained optimization
The rank-restricted problem (5) is nonconvex but lends itself well to more computationally efficient algorithms than (2). We consider two types of algorithm: a projected proximal gradient descent algorithm and the Burer-Monteiro factorized approach. We first present theoretical guarantees (Theorems 2 and 3) that are independent of Theorem 1 (assuming only the existence of a low-rank solution to (2)), then we combine these results with Theorem 1 to obtain specific guarantees in Corollary 1.
@@ 2.1 Projected spectral algorithm
Under matrix RIP, this algorithm (with suitable stepsize $\eta$) is guaranteed to converge to the low-rank global optimum:
@@
Let $\widehat{M}$ be a solution to (2), and suppose $r\geq\hat{r}\coloneqq\operatorname{rank}(\widehat{M})$. Suppose $\mathcal{A}$ has $(2r,\delta)$-RIP for some $\delta<1/3$. Then, for stepsize $\eta=\frac{1}{1+\delta}$ and any $M_{0}\in\mathbf{R}^{d_{1}\times d_{2}}$ with $\operatorname{rank}(M_{0})\leq r$, the iterates (6) satisfy
@@
Thus we have linear convergence to $\widehat{M}$. We prove this in Section 5 (it is the consequence of a somewhat more general result, Lemma 3, that may be of independent interest, though the result and its proof are straightforward adaptations of other works such as [7, Theorem 3]).
@@ 2.2 Burer-Monteiro factorization
Another popular solution approach to (5) is the smooth factored problem
$\min_{\begin{subarray}{c}X\in\mathbf{R}^{d_{1}\times r}\\ Y\in\mathbf{R}^{d_{2}\times r}\end{subarray}}~{}\frac{1}{2}\lVert y-\mathcal{A}(XY^{T})\rVert^{2}+\frac{\lambda}{2}(\lVert U\rVert_{\mathrm{F}}^{2}+\lVert V\rVert_{\mathrm{F}}^{2}),$ (7)
@@
Suppose $r\geq\hat{r}\coloneqq\operatorname{rank}(\widehat{M})$, where $\widehat{M}$ is a solution to (2), and suppose $\mathcal{A}$ has $(2r,\delta)$-RIP for some $\delta<1/3$. Then every second-order critical point $(X,Y)$ of (7) satisfies $XY^{T}=\widehat{M}$.
This follows from Theorem 2 by a minor adaptation (whose proof we omit) of [8, Theorem 2.3(a)]. Theorem 3 guarantees that local algorithms can reach the global optimum. For example, second-order methods such as trust-regions will reach second-order critical points (see, e.g., [9]), and gradient descent with random initialization and a suitable step size is known to converge to a local minimum [10]; Theorem 3 ensures that these limits will be global minima.
@@ 2.3 Combining the results
Then (2) has a unique solution $\widehat{M}$. Furthermore, choose
@@
and suppose $\mathcal{A}$ furthermore has $(2r,\delta)$-RIP for some $\delta<1/3$. Then
@@ 3.1 Estimator rank bounds in matrix sensing
Compared to our problem, it is much easier to bound estimator rank in noiseless matrix sensing because it is possible to recover the ground truth $M_{*}$ exactly. There is a plethora of results (see [11, 12] for surveys) showing that, for various conditions on the measurements $\mathcal{A}$ (including RIP) and various choices of low-rank–inducing penalty functions (including nuclear norm), if we observe $y=\mathcal{A}(M_{*})$ and solve a problem of the form
@@
the unique solution is $M_{*}$. Clearly the solution has low rank, but this entirely depends on the fact that we recover $M_{*}$ exactly. This can be extended to the noisy case when the error ($\xi$ in (1)) is sparse: that is, a large fraction of measurements remains uncorrupted. See, for example, [13, 14] for convex optimization estimators and [15, 16, 17, 18, 19] for the nonconvex case. For general non-sparse error, we cannot hope to recover $M_{*}$ exactly, so this approach does not work.
For general noisy matrix sensing, the closest result we are aware of to our Theorem 1 is [20, Theorem 8]. Similar results (with a nonconvex rank penalty rather than a nuclear norm) can be found in [21]. However, the estimators in these papers are not true least-squares estimators like (2) and do not yield consistency ($\widehat{M}\to M$) as the noise $\xi\to 0$. The analysis depends on the objective function being strongly convex on $\mathbf{R}^{d_{1}\times d_{2}}$, which is a far stronger condition than RIP on low-rank matrices.
@@ 3.2 Analogous results in sparse recovery
Many works in the LASSO literature show that the solution support (set of nonzero coordinates) is (approximately) the same (or a subset of) the ground truth support. A representative result and many further references can be found in [24, Chapter 11] (specifically, see Theorem 11.3). The analysis techniques depend on the fact that the support is a discrete quantity and therefore robust to some degree of noise. The low-rank matrix sensing problem has no equivalent discrete structure (the closest analog is the row and column subspaces, which are continuous), so we need a different analysis technique.
@@ 3.3 Low-rank optimization for matrix recovery
Many previous results consider algorithms of the form initialization + local optimization. These results show that, with a sufficiently accurate initial estimate (often obtained by a spectral method), a local algorithm (such as gradient descent or alternating minimization) will converge to a good solution. We do not attempt to cover the full body of such results (as they have a fundamentally different character than ours), but see the survey [25] as well as more recent papers such as [17, 26, 19, 27] for further references.
@@
where $f$ is a function whose global minimum is known to have low rank and whose Hessian has certain RIP-like properties. These results are very useful for deriving a nonconvex landscape result like Theorem 3 from a convex problem rank bound like Theorem 1. The paper [39] is particularly relevant because it explicitly considers the nuclear norm penalty and factored equivalent that we use.
Finally, there are many results [44, 45, 46, 47, 48, 49, 38] showing that, in the presence of (non-sparse) noise, we can still obtain a statistically benign landscape in the sense that every local minimum (even every second-order critical point) of (8) is close to the ground truth $M_{*}$ (where the error depends on the amount of noise). In the limit of zero noise, these results also yield exact recovery of the ground truth (and hence globally benign landscape).
@@
Importantly, none of the works mentioned above (initialization + convergence or global landscape) shows that we can efficiently obtain a global optimum to the nonconvex problem in the case of general non-sparse measurement noise. A notable counterpoint is [51], which shows that (8) can have a globally benign landscape even with noise. To our knowledge, this is the only such theoretical result prior to ours. The result has a very different character from ours; it depends fundamentally on the (Riemannian) geometry of the space of low-rank matrices. Hence it is only valid when we choose $r=\operatorname{rank}(M_{*})$ exactly, and it depends on the condition number $\frac{\sigma_{1}(M_{*})}{\sigma_{r}(M_{*})}$. Furthermore, it says nothing about solutions to (8) for larger values of $r$. Our result Theorem 3 exploits the connection between the nonconvex problem (7) and the convex problem (2), and thus we show that second-order critical points of (7) are globally optimal even without any rank restriction.
@@ 4.1 Preliminary facts
Many variations of this result can be found in the sparse and low-rank–matrix recovery literature (see, e.g., [3, proof of Theorem 3.3]). It can be proved by partitioning the singular values of $M$ into blocks of size $r$ in decreasing order; $M_{k}$ is the matrix formed by the singular values and vectors corresponding to the $k$th such block. Then, for $k\geq 2$, $\lVert M_{k}\rVert_{\mathrm{F}}\leq\frac{\lVert M_{k-1}\rVert_{*}}{\sqrt{r}}$.

### v2

@@ Abstract
We show that solutions to the popular convex matrix LASSO problem (nuclear-norm–penalized linear least-squares) have low rank under similar assumptions as required by classical low-rank matrix sensing error bounds. Although the purpose of the nuclear norm penalty is to promote low solution rank, a proof has not yet (to our knowledge) been provided outside very specific circumstances. Furthermore, we show that this result has significant theoretical consequences for nonconvex rank-constrained optimization approaches. Specifically, we show that if (a) the ground truth matrix has low rank, (b) the (linear) measurement operator has the matrix restricted isometry property (RIP), and (c) the measurement error is small enough relative to the nuclear norm penalty, then the LASSO solution is unique and has rank (approximately) bounded by that of the ground truth. From this, we show (a) that a low-rank–projected proximal gradient descent algorithm will converge linearly to the unique LASSO solution from any initialization, and (b) that the nonconvex landscape of the low-rank Burer-Monteiro–factored problem formulation is benign in the sense that all second-order critical points are globally optimal and yield the unique LASSO solution.
@@ 1 Introduction and main result
For such a problem, a popular choice of estimator is a solution to the nuclear-norm–regularized least-squares optimization problem, also know as the matrix LASSO:
@@
Curiously, although the entire point of the nuclear norm penalty in (2) is to promote low solution rank, and there are, as mentioned, many theoretical guarantees in the literature that LASSO solutions are good estimators of $M_{*}$ in the sense of error norm, there is little in the way of theoretical guarantees that the LASSO problem has a solution whose rank is, in fact, low (i.e., comparable to that of $M_{*}$). The precise rank of a solution is important for two practical reasons:
  • Storage and computation: The space required to store a solution $\widehat{M}$ with rank $\hat{r}$ is $O(\hat{r}(d_{1}+d_{2}))$. Similarly, the cost of multiplication by $\widehat{M}$ scales proportionally to $\hat{r}$. If $\hat{r}$ is large, we could try to save space and computational cost by taking a low-rank approximation of $\widehat{M}$, but this may introduce additional error.
  • Algorithmic efficiency: The optimization problem (2) is convex and has $d_{1}d_{2}$ variables. Although it can be solved in polynomial time, the computation may still be impractically expensive when $d_{1}$ and $d_{2}$ are large. If we knew that a low-rank solution existed, we could consider algorithms that only search over low-rank matrices; this is more practical for large problem sizes.
In this paper, we partially fill the theoretical gap for solution rank by showing that, under conditions already standard in the literature, the solution is unique and has rank similar to that of the ground truth $M_{*}$. Furthermore, we show that rank-resticted optimization algorithms (which scale better than a generic full-rank algorithm) can find this solution; see Section 2 for more details.
By arriving at nonconvex low-rank algorithms via the LASSO, we retain a key advantage of the LASSO: that it is rank-agnostic. We do not need to know $r^{*}$ precisely in order to obtain optimal error bounds. On the other hand, the existing theory for direct (unregularized) nonconvex approaches either assumes exact knowledge of $r^{*}$ or suffers increased sensitivity to noise if we overestimate the rank; see Section 3.3 for a literature review.
Our main solution rank result will depend on three properties of our measurements in (1):
@@
These concepts are well established in the matrix sensing literature. It is known [1, 2, 3] that, if $r^{*}=\operatorname{rank}(M_{*})$, and $\mathcal{A}$ has $(r,\delta)$-RIP for [^2] We write $a\lesssim b$ (equivalently, $b\gtrsim a$) to mean $a\leq Cb$ for some unspecified but universal constant $C>0$. We write $a\approx b$ to mean $a\lesssim b$ and $a\gtrsim b$ simultaneously. $r\gtrsim r^{*}$ and sufficiently small $\delta$, then any solution $\widehat{M}$ to (2) with regularization parameter $\lambda\approx\lVert\mathcal{A}^{*}(\xi)\rVert_{\ell_{2}}$ satisfies
$\lVert\widehat{M}-M_{*}\rVert_{\mathrm{F}}\lesssim\sqrt{r^{*}}\lVert\mathcal{A}^{*}(\xi)\rVert_{\ell_{2}}.$
This error is minimax-optimal (within constants) for Gaussian noise [1, 3]. This illustrates one key advantage of the matrix LASSO: we do not need to know the true rank $r^{*}$ precisely beforehand to obtain the optimal ($r^{*}$-dependent) error rate. In contrast, the error of unregularized low-rank methods (see Section 3.3) will, in general, suffer if the search rank is chosen too large.
We can now state our main result (proved in Section 4):
@@
The condition (4) is qualitatively similar to what was assumed for prior error bounds on $\widehat{M}$ (see above).
Note that there is some flexibility in how we choose $M_{*}$ and $\xi$. For example, if the ground truth is only approximately low-rank, we can decompose it into an exactly low-rank matrix plus a residual and then set $M_{*}$ to be the low-rank matrix and subsume the residual (through $\mathcal{A}$) into $\xi$. Again, see [1, 2] for examples of this type of analysis.
@@
The exact solution (which is unique due to strong convexity) can be obtained by singular value soft-thresholding $Y$: if $Y=U_{Y}\Sigma_{Y}V_{Y}^{T}$ is the singular value decomposition (SVD) of $Y$, then the solution is $\widehat{M}=U_{Y}(\Sigma_{Y}-\lambda I)_{+}V_{Y}^{T}$, where $(\Sigma_{Y}-\lambda I)_{+}$ replaces any negative diagonal elements of $\Sigma_{Y}-\lambda I$ by zero. If $\lVert\Xi\rVert_{\ell_{2}}\leq\lambda$, simple singular value perturbation arguments yield $\operatorname{rank}(\widehat{M})\leq\operatorname{rank}(M_{*})$. Our result Theorem 1 can be interpreted as extending this (modulo constants and the RIP constant $\delta^{*}$) to general RIP measurement operators $\mathcal{A}$.
@@ 2 Guarantees for nonconvex rank-constrained optimization
The rank-restricted problem (5) is nonconvex but lends itself well to more computationally efficient algorithms than (2). We consider two types of algorithm: a projected proximal gradient descent algorithm and the Burer-Monteiro factorized approach. We first present theoretical guarantees (Theorems 2 and 3) that are independent of Theorem 1, then we combine these results with Theorem 1 to obtain more specific results in Corollary 1.
@@ 2.1 Projected spectral algorithm
Under matrix RIP, if a low-rank global solution to (2) exists, this algorithm (with suitable stepsize $\eta$) is guaranteed to converge to it:
@@
Let $\widehat{M}$ be any solution to (2), and suppose $r\geq\hat{r}\coloneqq\operatorname{rank}(\widehat{M})$. Suppose $\mathcal{A}$ has $(2r,\delta)$-RIP for some $\delta<1/3$. Then, for stepsize $\eta=\frac{1}{1+\delta}$ and any $M_{0}\in\mathbf{R}^{d_{1}\times d_{2}}$ with $\operatorname{rank}(M_{0})\leq r$, the iterates (6) satisfy
@@
Thus we have linear convergence to $\widehat{M}$ (which must, in fact, be the only solution to (2) with rank at most $r$). We prove this in Section 5.1 (it is the consequence of a somewhat more general result, Lemma 3, that may be of independent interest, though the result and its proof are straightforward adaptations of other works such as [8, Theorem 3]).
@@ 2.2 Burer-Monteiro factorization
Another popular algorithmic approach to (5) is via the smooth factored problem
$\min_{\begin{subarray}{c}X\in\mathbf{R}^{d_{1}\times r}\\ Y\in\mathbf{R}^{d_{2}\times r}\end{subarray}}~{}\frac{1}{2}\lVert y-\mathcal{A}(XY^{T})\rVert^{2}+\frac{\lambda}{2}(\lVert X\rVert_{\mathrm{F}}^{2}+\lVert Y\rVert_{\mathrm{F}}^{2}),$ (7)
@@
Under the same conditions as Theorem 2, every second-order critical point $(X,Y)$ of (7) satisfies $XY^{T}=\widehat{M}$.
This follows from Theorem 2 by a straightforward adaptation of [9, Theorem 2.3(a)]. For completeness, we state and prove the adapted result as Lemma 4 in Section 5.2. A similar result to Theorem 3 (with a more direct proof) is [10, Thm. 1.7].
Theorem 3 guarantees that local algorithms can reach the global optimum. For example, second-order methods such as trust-regions will reach second-order critical points (see, e.g., [11]), and gradient descent with random initialization and a suitable step size is known to converge to a local minimum [12]; Theorem 3 ensures that these limits will be global minima.
@@ 2.3 Combining the results
Then (2) has a unique solution $\widehat{M}$. Furthermore, choose optimization rank parameter
@@
and suppose $\mathcal{A}$ has $(2r,\delta)$-RIP for some $\delta<1/3$. Then
@@ 3.1 Estimator rank bounds in matrix sensing
Compared to our problem, it is much easier to bound estimator rank in noiseless matrix sensing because it is possible to recover the ground truth $M_{*}$ exactly. There is a plethora of results (see [13, 14] for surveys) showing that, for various conditions on the measurements $\mathcal{A}$ (including RIP) and various choices of low-rank–inducing penalty functions (including nuclear norm), if we observe $y=\mathcal{A}(M_{*})$ and consider a problem of the form
@@
the only global optimum is $M_{*}$. Clearly, this has low rank, but this fact entirely depends on the fact that we recover $M_{*}$ exactly. This can be extended to the noisy case when the error ($\xi$ in (1)) is sparse: that is, a large fraction of measurements remains uncorrupted. See, for example, [15, 16] for convex optimization estimators and [17, 18, 19, 20, 21] for the nonconvex case. For general non-sparse error, we cannot hope to recover $M_{*}$ exactly, so this approach does not work.
For general noisy matrix sensing, the closest result we are aware of to our Theorem 1 is [22, Theorem 8]. Similar results (with a nonconvex rank penalty rather than a nuclear norm) can be found in [23]. However, the estimators in these papers are not true least-squares estimators like (2) and do not yield consistency ($\widehat{M}\to M_{*}$) as the noise $\xi\to 0$. The analysis depends on the objective function being strongly convex on $\mathbf{R}^{d_{1}\times d_{2}}$, which is a far stronger condition than RIP on low-rank matrices.
@@ 3.2 Analogous results in sparse recovery
Many works in the LASSO literature show that, under certain conditions, the support (set of nonzero coordinates) of any solution is (approximately) the same (or a subset of) the ground truth support. A representative result and many further references can be found in [26, Chapter 11] (specifically, see Theorem 11.3). The analysis techniques depend on the fact that the support is a discrete quantity and therefore robust to some degree of noise. The low-rank matrix sensing problem has no equivalent discrete structure (the closest analog is the row and column subspaces, which are continuous), so we need a different analysis technique.
@@ 3.3 Low-rank optimization for matrix recovery
Many previous results consider algorithms of the form initialization + local optimization. These results show that, with a sufficiently accurate initial estimate (often obtained by a spectral method), a local algorithm (such as gradient descent or alternating minimization) will converge to a global optimum or at least to a good estimate of the ground truth. We do not attempt to cover the full body of such results (as they have a fundamentally different character than ours), but see the survey [27] as well as more recent papers such as [19, 28, 21, 29] for further references. Of particular interest is [30], which proves error rates for Gaussian noise that are minimax optimal; however, this assumes precise knowledge of the ground truth rank $r^{*}$.
@@
where $f$ is a function whose global minimum is known to have low rank and whose Hessian has certain RIP-like properties. The paper [10] is particularly relevant because it explicitly considers the nuclear norm penalty and factored equivalent that we use; their Theorem 1.7 is very similar to our Theorem 3.
Finally, there are many results [46, 47, 48, 49, 50, 51, 41] showing that, in the presence of (non-sparse) noise, we can still obtain a statistically benign landscape in the sense that every local minimum (even every second-order critical point) of (8) is close to the ground truth $M_{*}$ (where the error depends on the amount of noise). In the limit of zero noise, these results also yield exact recovery of the ground truth (and hence globally benign landscape). However, none of these results, to our knowledge, guarantees optimal error rates with respect to noise as we find in, for example, [1, 2, 3] for the LASSO.
@@
Importantly, none of the works mentioned above (initialization + convergence or global landscape) shows that we can efficiently obtain a global optimum to the nonconvex problem in the case of general non-sparse measurement noise; these works contain guarantees that the resulting estimate is near the ground truth, but they leave open the possibility that there is a global optimum with strictly better objective function value. A notable exception is [53], which, among other things, shows that (8) can be free of non-global local minima even with (general) noise. To our knowledge, this is the only such theoretical result prior to ours. The result has a very different character from ours; it depends fundamentally on the (Riemannian) geometry of the space of low-rank matrices. Hence it is only valid when we choose $r=\operatorname{rank}(M_{*})$ exactly, and it depends on the condition number $\frac{\sigma_{1}(M_{*})}{\sigma_{r}(M_{*})}$. Furthermore, it says nothing about local or global optima of (8) for larger values of $r$. Our result Theorem 3 exploits the connection between the nonconvex problem (7) and the convex problem (2), and thus we show that second-order critical points of (7) are globally optimal even without any rank restriction.
@@ 4.1 Preliminary facts
Many variations of this result can be found in the sparse and low-rank–matrix recovery literature (see, e.g., [4, proof of Theorem 3.3]). It can be proved by partitioning the singular values of $M$ into blocks of size $r$ in decreasing order; $M_{k}$ is the matrix formed by the singular values and vectors corresponding to the $k$th such block. Then, for $k\geq 2$, $\lVert M_{k}\rVert_{\mathrm{F}}\leq\frac{\lVert M_{k-1}\rVert_{*}}{\sqrt{r}}$.